OpenStudy (anonymous):

If f(x) = |x + 7|, find f(–12).

OpenStudy (anonymous):

All you have to do is plug the -12 in for x and solve like you would any other equation like that.

OpenStudy (anonymous):

put x with -12 so it will be |-12+7|=|-5|=5

OpenStudy (campbell_st):

just substitute f(-12) = |-12 +7| with absolute value just find the convert your answer to the positive

OpenStudy (anonymous):

just substitute to x=-12 into the equation. Anything within the 'bar' will become positive, so |-5|=5.

OpenStudy (anonymous):

f(x) = |x + 7| means that f(x)=x+7, if x>=-7 and f(x)=-(x+7), if x<-7. Since -12<-7, f(x)=-x-7, so f(-12)=5.